OpenStudy (anonymous):

The length of a rectangle is 15 in and is dilated to 30 in. What is the scale factor?

OpenStudy (anonymous):

30=k *15

OpenStudy (anonymous):

scale factor=k

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

solve for k

OpenStudy (anonymous):

30=k *15 k=2 So, scale factor=2